OpenStudy (anonymous):

A family has two cars. The first car has a fuel efficiency of 40 miles per gallon of gas and the second has a fuel efficiency of 15 miles per gallon of gas. During one particular week, the two cars went a combined total of 1025 miles, for a total gas consumption of 35 gallons. How many gallons were consumed by each of the two cars that week?

OpenStudy (anonymous):

x = car with 40 mph gas y = car with 15 mph gas x + y = 35 --> x = 35 - y 40x + 15y = 1025 now sub 35 - y in for x in the 2nd equation 40x + 15y = 1025 40(35 - y) + 15y = 1025 1400 - 40y + 15y = 1025 1400 - 25y = 1025 -25y = 1025 - 1400 -25y = - 375 y = -375/-25 y = 15 now sub 15 in for y in the 1st equation x + y = 35 x + 15 = 35 x = 35 - 15 x = 20 check... 40x + 15y = 1025 40(20) + 15(15) = 1025 800 + 225 = 1025 1025 = 1025 the 40 mph (x) uses 20 gallons the 15 mph car uses 15 gallons

OpenStudy (anonymous):

any questions ?